# Mean-field Potts and random-cluster dynamics from high-entropy initializations

### v1

@@ 3.1 Quasi-equilibrating away from the unstable fixed point
Note that $\mu^{\mathsf{ord}}$ has $|\mathcal{L}_{1}(A)|$ concentrated around $\theta_{\mathsf{r}}>\theta_{*}$, where $\theta_{\mathsf{r}}$ is the (stable) fixed point of $f$ to the right of $\theta_{*}$. On the other hand $\mu^{\mathsf{dis}}$ has $|\mathcal{L}_{1}(A)|$ that is concentrated around 00.
@@
Suppose that $q>2$, $\gamma>0$ and $\beta\in(\beta_{\mathsf{u}},\beta_{\mathsf{s}})$. If $X_{0}$ is such that $|\mathcal{L}_{1}(X_{0})|\geq\theta_{*}n+\gamma\sqrt{n}$, $|\mathcal{L}_{2}(X_{0})|=O(\log n)$, and $R_{2}^{-}(X_{0})=O(n)$. There exists $C>0$ such that if $T=C\log n$, we have
@@
Let $q>2$, $\gamma>0$ and $\beta\in(\beta_{\mathsf{u}},\beta_{\mathsf{s}})$. If $X_{0}$ is any configuration such that $L_{0}\geq\theta_{*}n+\gamma\sqrt{n}$, $|\mathcal{L}_{2}(X_{0})|=O(\log n)$ and $R_{2}^{-}(X_{0})=O(n)$, then there exist constants $s\in(0,\theta_{r}-\theta_{*})$ and $C>0$ so that the first time that $L_{t}\geq(\theta_{*}+s)n$, $|\mathcal{L}_{2}({X}_{t})|=O(\log n)$, and $R_{2}^{-}({X}_{t})=O(n)$, all hold is at most $T=C\log n$ with probability $1-O(\gamma^{-2})$.
@@
Lemma 3.3 (Lemma 3.19 [2]).
Suppose $0<\beta<\beta_{\mathsf{s}}$ and that $X_{t}$ has at most one large component whose size is at least $2n^{11/12}$. Let $\varepsilon>0$ be a small constant. Then for sufficiently large $n$: if $L_{t}/n\geq\theta_{\mathsf{s}}+\varepsilon$, then
$\mathbb{E}[L_{t+1}\mid X_{t},\Lambda_{t+1}]\leq L_{t}+f(L_{t}/n)n+3n^{1/4},$
@@
Proof.
Let $\tau$ be the first time $t$ when $L_{t}\geq(\theta_{*}+s)n$ where $s>0$ is a small constant that will be decided. Let $(t_{i})_{i\geq 0}$ be the subset of times in $[T]=\{1,...,T\}$ at which the largest component is activated. We can condition on this sequence, generating $\sigma$-algebra $\mathscr{T}$, and notice that this can be generated by a sequence of $T$ independent $\text{Ber}(1/q)$ random variables which are also independent of the remaining randomness of the dynamics (say by reserving the first activation coin to always be used for the largest component).
Let $c_{*}=f^{\prime}(\theta_{*})/2>0$, and let $M$, $M_{0}$ and $M_{1}$ be sufficiently large constants. Our goal in this proof is to show the following by induction on $t$:
$\mathbb{P}\bigg{(}\bigcap_{k=1}^{t\wedge\tau}\big{\{}(L_{k}-\theta_{*}n)$ $\geq\Big{(}1+\frac{c_{*}}{4}\Big{)}^{\mathbf{1}\{\Lambda_{k}\}}(L_{k-1}-\theta_{*}n),R_{2}^{-}(X_{k})\leq Mn,|\mathcal{L}_{2}(X_{k})|\leq M_{1}n\big{\}}\mid\mathscr{T}\bigg{)}$
$\qquad\qquad\qquad\qquad\qquad\geq\prod_{i:t_{i}\leq t\wedge\tau}\Big{(}1-\frac{16M_{0}^{2}}{\gamma^{2}(1+\frac{c_{*}}{4})^{2i-2}c_{*}^{2}}\Big{)}\cdot\Big{(}1-O(T/n)\Big{)}\,.$ (14)
In words, (3.1) shows that with the stated probability, $L_{k}$ drifts away from $\theta_{*}n$ while preserving the control on $R_{2}^{-}(X_{k})$ and $|\mathcal{L}_{2}(X_{k})|$. Let us first conclude the proof assuming (3.1). For any given $s>0$, if $T=C\log n$ with sufficiently large $C>0$, then standard binomial tail bounds yield that |$|\{t_{i}:t_{i}\leq T\}|\geq T/2q$ with probability at least $1-n^{-1}$, and
@@
Thus, on any such realization of $\mathscr{T}$ where |$|\{t_{i}:t_{i}\leq T\}|\geq T/2q$, the event in (3.1) implies $\{\tau\leq T\}\cap\{R_{2}^{-}(X_{\tau})\leq Mn\}\cap\{|\mathcal{L}_{2}(X_{\tau})|\leq M_{1}n\}$. As such, we can conclude that
($\mathbb{P}\big{(}\{\tau\leq T\}$$\cap\{R_{2}^{-}(X_{\tau})\leq Mn\}\cap\{|\mathcal{L}_{2}(X_{\tau})|\leq M_{1}n\}\big{)}$
@@
We now proceed to prove (3.1) inductively. The event in the probability holds at initialization deterministically; now suppose (3.1) holds at time $t$ and let us show it holds at $t+1$. If $\tau=t$, then the event in the probability is unchanged at $t+1$, and the right-hand side is at least its value at $t$. Now suppose that $t<\tau$, and fix $(X_{k})_{k\leq t}$ such that the events in (3.1) hold for it.
If $t+1\notin\{t_{i}\}$ and $\Lambda_{t+1}^{c}$ occurs, then $L_{t+1}\geq L_{t}$ deterministically. In this case, it suffices to consider only the probability of $\{R_{2}^{-}(X_{t+1})\leq Mn,|\mathcal{L}_{2}(X_{t+1})|\leq M_{1}n\}$. On $\Lambda_{t+1}^{c}$, the decrease in $R_{2}^{-}$ as a result of the dissolution of active components is $R_{2}^{-}(X_{t})/q$ in expectation, and at least $R_{2}^{-}(X_{t})/q-o(n)$ with probability $1-O(n^{-1})$ by Hoeffding’s inequality and the fact that $R_{2}^{-}(X_{t})\leq Mn$ and $|\mathcal{L}_{2}(X_{t})|\leq M_{1}n$. Moreover, with probability $1-O(n^{-1})$, $A_{t+1}\in[\frac{n-L_{t}}{q}-\sqrt{n\log n},\frac{n-L_{t}}{q}+\sqrt{n\log n}]$. Thus, $A_{t}\cdot\frac{\beta}{n}<1$ is bounded away from $1$ uniformly in $n$. Lemma 2.8 implies that the increase in $R_{2}^{-}$ as a result of the creation of new components in the sub-critical percolation step is at most $C_{1}n$ with probability $1-O(n^{-1})$, for some constant $C_{1}>0$; besides, with probability $1-O(n^{-1})$, the sizes of new components in the sub-critical percolation step is at most $M_{1}\log n$. By a union bound,
$\mathbb{P}\big{(}|\mathcal{L}_{2}(X_{t+1})|\leq M_{1}n,R_{2}^{-}(X_{t+1})\leq R_{2}^{-}(X_{t})\big{(}1-\tfrac{1}{q}\big{)}+C_{1}n+o(n)\mid X_{t},\Lambda_{t+1}^{c}\big{)}=1-O(n^{-1}).$
@@
Now suppose $t+1=t_{i}$ for some $i\geq 1$ so that the events $\Lambda_{t+1}$, $R_{2}^{-}(X_{t})\leq Mn$, $|\mathcal{L}_{2}(X_{t})|\leq M_{1}n$ and
@@
($\mathbb{P}(\{R_{2}^{-}(X_{t+1})\leq Mn,|\mathcal{L}_{2}(X_{t+1})|\leq M_{1}n\}\mid X_{t},\Lambda_{t+1})=1-O(n^{-1}).$ (16)
@@
$f\big{(}\frac{L_{t}}{n}\big{)}n-2n^{\frac{1}{4}}-s_{t}\geq f\big{(}\frac{L_{t}}{n}\big{)}\cdot\frac{n}{2}.$ (19)
@@
where the last inequality follows from the Fact 3.26 that there exist constants $M_{0}>0$ and $s\in(0,1)$ such that if $\theta_{*}n-sn\leq L_{t}\leq\theta_{*}n+sn$, then
@@
$\mathbb{P}\big{(}L_{t+1}-\theta_{*}n\geq(1+\frac{c_{*}}{4})\cdot(L_{t}-\theta_{*}n)\mid X_{t},\Lambda_{t+1}\big{)}$ $\geq\mathbb{P}\big{(}L_{t+1}\geq L_{t}+f(L_{t}/n)\cdot\frac{n}{2}\mid X_{t},\Lambda_{t+1}\big{)}$
@@
By a union bound, this inequality and (16) conclude the induction step of (3.1) for the case of $\Lambda_{t+1}$. ∎
@@
Now, let $\Omega_{\mathsf{good}}$ be the set of all random-cluster configurations $X$ such that $|\mathcal{L}_{1}(X)|\geq(\theta_{*}+s)n$, $|\mathcal{L}_{2}({X})|=O(\log n)$ and $R_{2}^{-}({X})=O(n)$. Lemma 3.25 from [2] shows that if $X_{t}\in\Omega_{\mathsf{good}}$, then $X_{t+1}\in\Omega_{\mathsf{good}}$ with probability $1-O(n^{-1})$. (Again, this result from [2] is stated for $\beta\geq q$, but it extends to the $q>2$ and $\beta>\beta_{\mathsf{u}}$ setting by the same observations made above about the subcriticallity/supercriticality of the percolation step; in addition, Lemma 3.25 only states a $1-o(1)$ but its proofs yields a $1-O(n^{-1})$ bound on the probability.) By averaging over all the configurations on $\Omega_{\mathsf{good}}$, we get from (21) that
@@
Since $\Delta_{0}=O(n)$, there exists some $T=O(\log n)$ so that $\mathbb{E}[\Delta_{T}]\leq C\sqrt{n}$, and by Markov’s inequality we have $\Delta_{T}\leq 2C\sqrt{n}/\varepsilon$ with probability $21-\varepsilon/$ 2 for any fixed $\varepsilon>0$. Finally, note that $X_{T}\in\Omega_{\mathsf{good}}$ with probability $1-o(1)$ and in the percolation step of the last step, Lemma 2.8 and 2.11 imply that $I_{1}(X_{T})=\Omega(n)$ with probability $1-o(1)$. The result then follows from a union bound. ∎
@@
Our first observation is that Lemma 3.27 and Corollary 3.33 from [2] hold when $q>2$ and $\beta>\beta_{\mathsf{u}}$. (This is a byproduct of the percolation step of the CM dynamics being subcritical (resp. supercritical) when the largest component of the configuration is inactive (resp., active).) To boost the probability of success of the coupling, we note that at a suitable $T_{0}=O(\log n)$, either the coupling to the same component structure succeeds with probability at least $\alpha=\Omega(1)$, or, by Corollary 3.33 from [2], both $X_{T_{0}}$ and $Y_{T_{0}}$ satisfy conditions 2, 3 and 4 from the lemma statement with probability $1-o(1)$ and also $||\mathcal{L}_{1}(X_{T_{0}})|-\theta_{\mathsf{r}}n|=O(\sqrt{n}\log^{2}n)$, $||\mathcal{L}_{1}(Y_{T_{0}})|-\theta_{\mathsf{r}}n|=O(\sqrt{n}\log^{2}n)$. Then, by Lemma 3.3, at time $T_{1}=T_{0}+O(\log n)$, we have that all four conditions from the lemma statement hold with probability $1-1/A$ for any desired constant $A>0$. Iterating this reasoning, we obtain a coupling for which ${X}_{kT_{1}}={Y}_{kT_{1}}$ with probability at least $1-(1-\alpha)^{k}-k/A-o(1)$ for any constant $k$. Letting $T=kT_{1}$ with $k$ and $A$ sufficiently large, we obtain a coupling under which $X_{T}$ and $Y_{T}$ have the same component structure with probability at least $21-\varepsilon/$ 2 for any $\varepsilon>0$. The result then follows from Lemma 3.16 in [2] and a union bound. ∎
@@
Let $q>2$, $s>0$ and $\beta\in(\beta_{\mathsf{u}},\beta_{\mathsf{s}})$. If $X_{0}$ is a configuration such that $L_{0}\leq(\theta_{*}-s)n$, $|\mathcal{L}_{2}(X_{0})|=O(\log n)$ and $R_{2}^{-}(X_{0})=O(n)$, then there exists $T=O(\log n)$ such that $L_{T}=O(\log n)$ and $I_{1}(X_{T})=\Omega(n)$ with probability $1-o(1)$.
@@
$\tau_{\varepsilon}:=\min\{t>0:L_{t}/n\notin[\theta_{\mathsf{s}}-\varepsilon,\theta_{*}-s/2]\}\qquad\text{and}\qquad\tau_{2}=\min\{t>0:|\mathcal{L}_{2}(X_{t})|>2n^{11/12}\}\,.$
@@
Suppose $(X_{t})_{t\geq 0}$ is initialized from $X_{0}$ with $L_{0}\geq\theta_{*}n+\gamma\sqrt{n}$, $|\mathcal{L}_{2}(X_{0})|=O(\log n)$, and $R_{2}^{-}(X_{0})=O(n)$. Let $(Y_{t})_{t\geq 0}$ be initialized from $\mu^{\mathsf{ord}}$ and restricted to $\Omega^{\mathsf{ord}}$ by rejecting any update that would take it out of $\Omega^{\mathsf{ord}}$; by Lemma 2.1 and a union bound, with probability $1-o(1)$, $(Y_{t})_{t\geq 0}$ does not feel the restriction to $\Omega^{\mathsf{ord}}$ (i.e., doesn’t attempt to leave $\Omega^{\mathsf{ord}}$) for exponential in $n$ many steps, so for all $e^{o(n)}$ time steps, we can treat $Y_{t}$ as the (unrestricted) CM chain initialized from $\mu^{\mathsf{ord}}$. We claim that there is $T=O(\log n)$ such that with probability $1-O(\gamma^{-2})$, we have coupled $X_{T}=Y_{T}$, which would imply the claim. In order to see this, notice that Lemma 3.2 can be stitched with Lemma 3.4 to get, with probability $1-O(\gamma^{-2})$, a configuration $X_{T_{1}}$ on which Lemma 3.5 can be applied. At the same time, $Y_{T_{1}}$ satisfies the conditions of Lemma 3.5, being a sample from $\mu^{\mathsf{ord}}$ (which satisfies those conditions by Lemma 2.1).

### v2

@@ 3.1 Quasi-equilibrating away from the unstable fixed point
Note that $\mu^{\mathsf{ord}}$ has $|\mathcal{L}_{1}(A)|$ concentrated around $\theta_{\mathsf{r}}>\theta_{*}$, where $\theta_{\mathsf{r}}$ is the (stable) fixed point of $f$ to the right of $\theta_{*}$. On the other hand $\mu^{\mathsf{dis}}$ has $|\mathcal{L}_{1}(A)|$ that is concentrated around 0.
@@
Suppose that $q>2$, $\gamma>0$ and $\beta\in(\beta_{\mathsf{u}},\beta_{\mathsf{s}})$. If $X_{0}$ is such that $|\mathcal{L}_{1}(X_{0})|\geq\theta_{*}n+\gamma\sqrt{n}$, $|\mathcal{L}_{2}(X_{0})|=O(\log n)$, and $R_{2}^{-}(X_{0})=O(n)$, then there exists $C>0$ such that if $T=C\log n$, we have
@@
Let $q>2$, $\gamma>0$ and $\beta\in(\beta_{\mathsf{u}},\beta_{\mathsf{s}})$. Let $X_{0}$ be any configuration such that $L_{0}\geq\theta_{*}n+\gamma\sqrt{n}$, $|\mathcal{L}_{2}(X_{0})|=O(\log n)$ and $R_{2}^{-}(X_{0})=O(n)$. Let $T_{s}$ be the first $t$ such that $L_{t}\geq(\theta_{*}+s)n$, $|\mathcal{L}_{2}({X}_{t})|=O(\log n)$, and $R_{2}^{-}({X}_{t})=O(n)$ all hold. Then, there exist constants $s\in(0,\theta_{r}-\theta_{*})$ and $C>0$ such that $T_{s}\leq C\log n$ with probability $1-O(\gamma^{-2})$.
@@
Lemma 3.3 (Lemma 3.19 and (3.11) of its proof, [2]).
Suppose $0<\beta<\beta_{\mathsf{s}}$ and that $X_{t}$ has at most one large component whose size is at least $2n^{11/12}$. Let $\varepsilon>0$ be a small constant. If $L_{t}/n\geq\theta_{\mathsf{s}}+\varepsilon$, then
$L_{t}+f(L_{t}/n)n-3n^{1/4}\leq\mathbb{E}[L_{t+1}\mid X_{t},\Lambda_{t+1}]\leq L_{t}+f(L_{t}/n)n+3n^{1/4},$
@@
We will now move on to prove Lemma 3.2.
Proof of Lemma 3.2.
Let $\tau$ be the first time $t$ when $L_{t}\geq(\theta_{*}+s)n$ where $s>0$ is a small constant that will be decided. Let $(t_{i})_{i\geq 0}$ be the subset of times in $T]=\{1,...,T\}$ at which the largest component is activated. We can condition on this sequence, generating $\sigma$-algebra $\mathscr{T}$, and notice that this can be generated by a sequence of $T$ independent $\text{Ber}(1/q)$ random variables which are also independent of the remaining randomness of the dynamics (say by reserving the first activation coin to always be used for the largest component).
Let $c_{*}=f^{\prime}(\theta_{*})/2>0$, and let $M$, $M_{0}$ and $M_{1}$ be sufficiently large constants. Consider the event:
$\mathcal{A}_{k}=\big{\{}(L_{k}-\theta_{*}n)\geq\Big{(}1+\frac{c_{*}}{4}\Big{)}^{\mathbf{1}\{\Lambda_{k}\}}(L_{k-1}-\theta_{*}n),R_{2}^{-}(X_{k})\leq Mn,|\mathcal{L}_{2}(X_{k})|\leq M_{1}\log n\big{\}}.$
Our goal in this proof is to show the following by induction on $t$:
$\mathbb{P}\bigg{(}\bigcap_{k=1}^{t\wedge\tau}\mathcal{A}_{k}\mid\mathscr{T}\bigg{)}\geq\prod_{i:t_{i}\leq t\wedge\tau}\Big{(}1-\frac{16M_{0}^{2}}{\gamma^{2}(1+\frac{c_{*}}{4})^{2i-2}c_{*}^{2}}\Big{)}\cdot\Big{(}1-O(T/n)\Big{)}\,.$ (14)
In words, (14) shows that with the stated probability, $L_{k}$ drifts away from $\theta_{*}n$ while preserving the control on $R_{2}^{-}(X_{k})$ and $|\mathcal{L}_{2}(X_{k})|$. Let us first conclude the proof assuming (14). For any given $s>0$, if $T=C\log n$ with sufficiently large $C>0$, then standard binomial tail bounds yield that $|\{t_{i}:t_{i}\leq T\}|\geq T/2q$ with probability at least $1-n^{-1}$, and
@@
Thus, on any such realization of $\mathscr{T}$ where $|\{t_{i}:t_{i}\leq T\}|\geq T/2q$, the event in (14) implies $\{\tau\leq T\}\cap\{R_{2}^{-}(X_{\tau})\leq Mn\}\cap\{|\mathcal{L}_{2}(X_{\tau})|\leq M_{1}\log n\}$. As such, we can conclude that
$\mathbb{P}\big{(}\{\tau\leq T\}$ $\cap\{R_{2}^{-}(X_{\tau})\leq Mn\}\cap\{|\mathcal{L}_{2}(X_{\tau})|\leq M_{1}\log n\}\big{)}$
@@
We now proceed to prove (14) inductively. The event in the probability holds at initialization deterministically; now suppose (14) holds at time $t$ and let us show it holds at $t+1$. If $\tau=t$, then the event in the probability is unchanged at $t+1$, and the right-hand side is at least its value at $t$. Now suppose that $t<\tau$, and fix $(X_{k})_{k\leq t}$ such that the events in (14) hold for it.
If $t+1\notin\{t_{i}\}$ and $\Lambda_{t+1}^{c}$ occurs, then $L_{t+1}\geq L_{t}$ deterministically. In this case, it suffices to consider only the probability of $\{R_{2}^{-}(X_{t+1})\leq Mn,|\mathcal{L}_{2}(X_{t+1})|\leq M_{1}\log n\}$. On $\Lambda_{t+1}^{c}$, the decrease in $R_{2}^{-}$ as a result of the dissolution of active components is $R_{2}^{-}(X_{t})/q$ in expectation, and at least $R_{2}^{-}(X_{t})/q-o(n)$ with probability $1-O(n^{-1})$ by Hoeffding’s inequality and the fact that $R_{2}^{-}(X_{t})\leq Mn$ and $|\mathcal{L}_{2}(X_{t})|\leq M_{1}\log n$. Moreover, with probability $1-O(n^{-1})$, $A_{t+1}\in[\frac{n-L_{t}}{q}-\sqrt{n\log n},\frac{n-L_{t}}{q}+\sqrt{n\log n}]$. Thus, $A_{t}\cdot\frac{\beta}{n}<1$ is bounded away from $1$ uniformly in $n$. Lemma 2.8 implies that the increase in $R_{2}^{-}$ as a result of the creation of new components in the sub-critical percolation step is at most $C_{1}n$ with probability $1-O(n^{-1})$, for some constant $C_{1}>0$; besides, with probability $1-O(n^{-1})$, the sizes of new components in the sub-critical percolation step is at most $M_{1}\log n$. By a union bound,
$\mathbb{P}\big{(}$ $|\mathcal{L}_{2}(X_{t+1})|\leq M_{1}\log n,R_{2}^{-}(X_{t+1})\leq R_{2}^{-}(X_{t})\big{(}1-\tfrac{1}{q}\big{)}+C_{1}n+o(n)\mid X_{t},\Lambda_{t+1}^{c}\big{)}$
$=1-O(n^{-1}).$
@@
Now suppose $t+1=t_{i}$ for some $i\geq 1$ so that the events $\Lambda_{t+1}$, $R_{2}^{-}(X_{t})\leq Mn$, $|\mathcal{L}_{2}(X_{t})|\leq M_{1}\log n$ and
@@
$\mathbb{P}(\{R_{2}^{-}(X_{t+1})\leq Mn,|\mathcal{L}_{2}(X_{t+1})|\leq M_{1}\log n\}\mid X_{t},\Lambda_{t+1})=1-O(n^{-1}).$ (16)
@@
$f\big{(}\frac{L_{t}}{n}\big{)}n-3n^{\frac{1}{4}}-s_{t}\geq f\big{(}\frac{L_{t}}{n}\big{)}\cdot\frac{n}{2}.$ (19)
@@
where the last equality follows from the Fact 3.26 that there exist constants $M_{0}>0$ and $s\in(0,1)$ such that if $\theta_{*}n-sn\leq L_{t}\leq\theta_{*}n+sn$, then
@@
$\mathbb{P}\big{(}L_{t+1}-\theta_{*}n\geq(1+\frac{c_{*}}{4})$ $\cdot(L_{t}-\theta_{*}n)\mid X_{t},\Lambda_{t+1}\big{)}$
$\geq\mathbb{P}\big{(}L_{t+1}\geq L_{t}+f(L_{t}/n)\cdot\frac{n}{2}\mid X_{t},\Lambda_{t+1}\big{)}$
@@
By a union bound, this inequality and (16) conclude the induction step of (14) for the case of $\Lambda_{t+1}$. ∎
@@
Now, let $\Omega_{\mathsf{good}}$ be the set of all random-cluster configurations $X$ such that $|\mathcal{L}_{1}(X)|\geq(\theta_{*}+s)n$, $|\mathcal{L}_{2}({X})|=O(\log n)$ and $R_{2}^{-}({X})=O(n)$. Lemma 3.25 from [2] shows that if $X_{t}\in\Omega_{\mathsf{good}}$, then $X_{t+1}\in\Omega_{\mathsf{good}}$ with probability $1-O(n^{-1})$. (Again, this result from [2] is stated for $\beta\geq q$, but it extends to the $q>2$ and $\beta>\beta_{\mathsf{u}}$ setting by the same observations made above about the subcriticallity/supercriticality of the percolation step; in addition, Lemma 3.25 only states a 1o(1)1-o(1) but its proofs yields a $1-O(n^{-1})$ bound on the probability.) By averaging over all the configurations on $\Omega_{\mathsf{good}}$, we get from (21) that
@@
Since $\Delta_{0}=O(n)$, there exists some $T=O(\log n)$ so that $\mathbb{E}[\Delta_{T}]\leq C\sqrt{n}$, and by Markov’s inequality we have $\Delta_{T}\leq 2C\sqrt{n}/\varepsilon$ with probability $1-\varepsilon/2$ for any fixed $\varepsilon>0$. Finally, note that $X_{T}\in\Omega_{\mathsf{good}}$ with probability 1o(1)1-o(1) and in the percolation step of the last step, Lemma 2.8 and 2.11 imply that $I_{1}(X_{T})=\Omega(n)$ with probability 1o(1)1-o(1). The result then follows from a union bound. ∎
@@
Our first observation is that Lemma 3.27 and Corollary 3.33 from [2] hold when $q>2$ and $\beta>\beta_{\mathsf{u}}$. (This is a byproduct of the percolation step of the CM dynamics being subcritical (resp. supercritical) when the largest component of the configuration is inactive (resp., active).) To boost the probability of success of the coupling, we note that at a suitable $T_{0}=O(\log n)$, either the coupling to the same component structure succeeds with probability at least $\alpha=\Omega(1)$, or, by Corollary 3.33 from [2], both $X_{T_{0}}$ and $Y_{T_{0}}$ satisfy conditions 2, 3 and 4 from the lemma statement with probability 1o(1)1-o(1) and also $||\mathcal{L}_{1}(X_{T_{0}})|-\theta_{\mathsf{r}}n|=O(\sqrt{n}\log^{2}n)$, $||\mathcal{L}_{1}(Y_{T_{0}})|-\theta_{\mathsf{r}}n|=O(\sqrt{n}\log^{2}n)$. Then, by Lemma 3.4, at time $T_{1}=T_{0}+O(\log n)$, we have that all four conditions from the lemma statement hold with probability 11/A1-1/A for any desired constant $A>0$. Iterating this reasoning, we obtain a coupling for which ${X}_{kT_{1}}={Y}_{kT_{1}}$ with probability at least $1-(1-\alpha)^{k}-k/A-o(1)$ for any constant $k$. Letting $T=kT_{1}$ with $k$ and $A$ sufficiently large, we obtain a coupling under which $X_{T}$ and $Y_{T}$ have the same component structure with probability at least $1-\varepsilon/2$ for any $\varepsilon>0$. The result then follows from Lemma 3.16 in [2] and a union bound. ∎
@@
Let $q>2$, $s>0$ and $\beta\in(\beta_{\mathsf{u}},\beta_{\mathsf{s}})$. If $X_{0}$ is a configuration such that $L_{0}\leq(\theta_{*}-s)n$, $|\mathcal{L}_{2}(X_{0})|=O(\log n)$ and $R_{2}^{-}(X_{0})=O(n)$, then there exists $T=O(\log n)$ such that $L_{T}=O(\log n)$ and $I_{1}(X_{T})=\Omega(n)$ with probability 1o(1)1-o(1).
@@
$\tau_{\varepsilon}:=\min\{t>0:L_{t}/n\notin[\theta_{\mathsf{s}}-\varepsilon,\theta_{*}-s/2]\}\quad\text{and}\quad\tau_{2}=\min\{t>0:|\mathcal{L}_{2}(X_{t})|>2n^{11/12}\}\,.$
@@
Suppose $(X_{t})_{t\geq 0}$ is initialized from $X_{0}$ with $L_{0}\geq\theta_{*}n+\gamma\sqrt{n}$, $|\mathcal{L}_{2}(X_{0})|=O(\log n)$, and $R_{2}^{-}(X_{0})=O(n)$. Let $(Y_{t})_{t\geq 0}$ be initialized from $\mu^{\mathsf{ord}}$ and restricted to $\Omega^{\mathsf{ord}}$ by rejecting any update that would take it out of $\Omega^{\mathsf{ord}}$; by Lemma 2.1 and a union bound, with probability 1o(1)1-o(1), $(Y_{t})_{t\geq 0}$ does not feel the restriction to $\Omega^{\mathsf{ord}}$ (i.e., doesn’t attempt to leave $\Omega^{\mathsf{ord}}$) for exponential in $n$ many steps, so for all $e^{o(n)}$ time steps, we can treat $Y_{t}$ as the (unrestricted) CM chain initialized from $\mu^{\mathsf{ord}}$. We claim that there is $T=O(\log n)$ such that with probability $1-O(\gamma^{-2})$, we have coupled $X_{T}=Y_{T}$, which would imply the claim. In order to see this, notice that Lemma 3.2 can be stitched with Lemma 3.4 to get, with probability $1-O(\gamma^{-2})$, a configuration $X_{T_{1}}$ on which Lemma 3.5 can be applied. At the same time, $Y_{T_{1}}$ satisfies the conditions of Lemma 3.5, being a sample from $\mu^{\mathsf{ord}}$ (which satisfies those conditions by Lemma 2.1).